an object of mass 2kg moving with a velocity of 3ms collides head on with another object of mass 1kg moving in the opposite direction with a velocity of 4ms the object stick together after collision,calculate their common speed
1
Expert's answer
2018-03-14T12:28:08-0400
The conservation law gives momentum
m_1 v_1-m_2 v_2=(m_1 〖+m〗_2 )u
u=(m_1 v_1-m_2 v_2)/(m_1 〖+m〗_2 )=(2×3-1×4)/(2+1)=2/3=0.67 m/s